# Personality Structured Interview for Large Language Model Simulation in Personality Research

### v1

@@ Abstract
Although psychometrics researchers have recently explored the use of large language models (LLMs) as proxies for human participants, LLMs often fail to generate heterogeneous data with human-like diversity, which diminishes their value in advancing social science research. To address these challenges, we explored the potential of the theory-informed Personality Structured Interview (PSI) as a tool for simulating human responses in personality research. In this approach, the simulation is grounded in nuanced real-human interview transcripts that target the personality construct of interest. We have provided a growing set of 357 structured interview transcripts from a representative sample [^1] Dataset and code will be shared later., each containing an individual’s response to 32 open-ended questions carefully designed to gather theory-based personality evidence. Additionally, grounded in psychometric research, we have summarized an evaluation framework to systematically validate LLM-generated psychometric data. Results from three experiments demonstrate that well-designed structured interviews could improve human-like heterogeneity in LLM-simulated personality data and predict personality-related behavioral outcomes (i.e., organizational citizenship behaviors and counterproductive work behavior). We further discuss the role of theory-informed structured interviews in LLM-based simulation and outline a general framework for designing structured interviews to simulate human-like data for psychometric research.
Personality Structured Interview for Large Language Model Simulation in Personality Research
Pengda Wang1, Huiqi Zou3, Hanjie Chen2, Tianjun Sun1, Ziang Xiao3, and Frederick L. Oswald1 1Department of Psychological Sciences, Rice University 2Department of Computer Science, Rice University 3Department of Computer Science, Johns Hopkins University {pw32,hc86,ts110,foswald}@rice.edu; {hzou11,ziang.xiao}@jhu.edu
@@ 1 Introduction
Personality research is an important field in psychology for understanding how individual traits shape significant life outcomes and trajectories, including career success, mental health, and overall well-being (e.g., Judge et al., 2002; Roberts et al., 2007; Robins et al., 2002). For instance, measures of conscientiousness have been consistently and positively correlated with job performance and academic success (Barrick and Mount, 1991), whereas measures of neuroticism are found to be strongly and positively correlated with mental health issues, such as anxiety and depression (John et al., 2008).
However, studying personality requires understanding habits of thoughts, attitudes, and behaviors across diverse contexts and cultures, which is often a research challenge. For example, some research questions involve rare or extreme cases, making it difficult to gather sufficient samples (e.g., Lynam and Widiger, 2001). Others require longitudinal designs to track personality development or dynamic social interactions (e.g., Damian et al., 2019; Roberts et al., 2006). Additionally, certain experimental manipulations, like inducing stress or altering life circumstances, may be feasible or unethical for causal insights Fisher (2021). If a method could accurately simulate the human personality distributions, they could supplement and accelerate personality research by offering a scalable, cost-effective approach to supplement traditional data collection and experimentation (Messeri and Crockett, 2024).
Several prior studies have supported this possibility, showing that LLMs can generate responses for personality scales that reflect personality traits resembling those of humans (e.g., Lee et al. 2024, Huang et al. 2023). However, studies found such methods face significant limitations in capturing individual differences at the item level (e.g., Wang et al. 2024a). Information at the item level provides a more granular understanding of how traits manifest in specific individuals and situations rather than relying solely on broad trait averages. However, without additional empirical information, items confound these sources of variance. Moreover, key psychometric challenges remain in modeling LLM responses, as they may not accurately reflect the intended personality construct.
To address this gap, we explored the potential of the Personality Structured Interview (PSI), which employs theory-informed questions to elicit information embedded in human narrative responses that are relevant to the target personality constructs of interest in LLM simulations. We believe this sophisticated linguistic approach can inform and enhance the measurement of human-like traits in heterogeneous LLM-simulated personality data, potentially leading to a greater understanding of personality.
Grounded in psychometric research (e.g., Cronbach, 1951; Cronbach and Meehl, 1955), we have summarized a framework for evaluating psychometric data that fully considers its hierarchical structure, in which observed responses (item level) are mapped to latent traits (domain level). This evaluation framework encompasses a range of analyses, from overall descriptive statistics to a more in-depth examination of psychometric performance.
Based on our evaluation framework, we conducted three experiments. Those three experiments evaluated our method’s effectiveness in replicating individual personality traits, simulating human-like personality distribution, and demonstrating personality-related behaviors with respect to established personality theory. We found our PSI method improves human-like diversity in personality traits simulation and recovers nuances in personality-related behaviors.
@@
Figure 1: Overview of the Development of the Personality Structured Interview and Experimental Implementation in This Study
Note: For an overview of the personality structured interview development process, refer to §3 and Appendix B. The structured interview development framework is detailed in Appendix C. Information on data collection and the human sample is provided in §4 and Appendix D. The LLM simulation process is described in §4 and Appendix F. Details on the experimental setup and results of the three experiments can be found in §4, §5, and Appendix I. For the psychometric data evaluation framework, see Appendix H.
In summary, our paper offers three contributions:
  • A theory-informed LLM-based simulation method (PSI) for personality research and a development framework (see Appendix C).
  • A growing dataset, now with 357 structured interview transcripts.
  • An evaluation framework for LLM-simulated psychometric data that is grounded in psychological theories (see Appendix H).
We further discuss the potential of PSI methods in advancing research (see Appendix C). We believe the framework behind PSI could be generalized to other psychological constructs and other psychometric data to develop theory-informed simulation methods for human understanding.
2 Personality Research and Related Works
Here we provide a concise overview of key personality concepts and discuss related works.
@@ 2 Personality Research and Related Works
MacKinnon (1944) proposed two definitions of personality. One emphasizes internal factors like temperament and interpersonal strategies that drive consistent behavior across time, situations, and cultures. The other focuses on interpersonal characteristics as perceived by others, linking personality to reputation. The former highlights internal drives, while the latter centers on external behaviors and social perception. Together, they underscore personality’s role in shaping thought patterns and behaviors in social interactions (Hogan et al., 1996).
@@
Personality encodes rich and complex information in language and text (Goldberg, 1990; Saucier and Goldberg, 2001). In fact, the Five-Factor Model (FFM) of personality is extensively researched (e.g., Costa and McCrae, 2008; John, 1999; McCrae and Costa Jr, 1997); it is directly based on the lexical hypothesis, which posits that individual differences that are important in human interactions (e.g., have survival value across cultures) are often encoded in some or all languages of the world. The five factors are openness, conscientiousness, extraversion, agreeableness, and neuroticism (OCEAN). From Galton’s (1884) early research to Allport and Odbert’s (1936) systematic organization and factor analysis of lexical terms, and through further developments by Norman (1963) and Goldberg (1990), the FFM theory has progressively evolved (see Appendix A for further details on the structure of personality). In other words, because language and text contain rich and complex information about personality, LLMs may capture and model such encoding by learning from vast amounts of training data.
There are many studies to-date on LLMs that focus on personality, where researchers aim to benchmark their psychological profiles (Lee et al., 2024; Li et al., 2024; Pellert et al., 2023), investigate their ability to simulate personality (Huang et al., 2023; Jiang et al., 2023; Serapio-García et al., 2023), and examine whether LLM agents align with character personality settings and how these settings influence their generated dialogues (Shao et al., 2023; Wang et al., 2024b; Xu et al., 2024).
@@
Currently, common methods for LLM simulation of human personality distributions have several limitations. The Persona-Chat dataset (Persona method) by Zhang et al. (2018) was originally designed to enhance the engagement of chit-chat models by increasing personalization. Therefore, its focus is more on enhancing personalization, rather than capturing personality traits. The adjective-based categorization method (Shape method) proposed by Serapio-García et al. (2023) effectively simulates profile-specific patterns of high/low standing on personality dimensions. However, this method presents a challenge in simulating human personality distributions, as restricting each case to a single dimension fails to capture the inherently multidimensional and hierarchical nature of personality data (e.g., Kachur et al., 2020).
To address the aforementioned limitations of personality-simulated data, we explore the use of theory-informed PSI transcription to simulate personality data.
3 Personality Structured Interview
The theory-informed structured interview mentioned in this study differs from previous interview approaches used in LLM-based simulations, such as the Life Interview  (Park et al., 2024), which broadly inquires about various aspects of an individual’s life. In contrast, the theory-informed structured interview is specifically designed to target the personality constructs that we aim to model.
A theory-informed approach is essential not only for improving the accuracy of data simulation but also for enhancing interpretability. Our goal is to generate data that not only aligns with actual data at the personality scale level but also exhibits theoretically coherent personality-related behaviors. By grounding LLM-generated data in theory, we ensure that the simulation goes beyond reproducing statistical properties to capturing meaningful psychological patterns. This enhances interpretability, facilitates transparent comparisons with human responses, and preserves the psychological validity of the constructs.
Learning from psychometric theories, we summarized a framework for developing structured interviews, which can serve as a template for creating interviews designed to collect targeted information for simulating specific types of psychometric data (see Appendix C). We present a PSI example designed to capture personality-related information, with relevant personality theories and question design detailed in §3.1.
We want to further emphasize that the “structured interview questions” mentioned here differ from those described by Wang et al. (2024b); Ran et al. (2024), who refers to converting personality scale items into open-ended questions. The main issue with their approach is that, while it attempts to rewrite items as open-ended questions to elicit richer responses, it may fail to achieve the intended effect. For instance, in their example, “Values artistic, aesthetic experiences” is rewritten as “Do you value artistic, aesthetic experiences?” Although this question appears open-ended, it tends to prompt a simple “yes” or “no” response and function as if it is close-ended, limiting the exploration of the respondent’s deeper thoughts or reasoning (Trull et al., 1998).
Moreover, such questions differ from the behavioral or situational questions typically found in structured interviews. Structured interviews usually involve more specific scenarios or tasks that encourage respondents to describe past experiences or hypothetical reactions (Campion et al., 1997). For example, to explore a respondent’s views on artistic and aesthetic experiences, a more effective question might be, “Please describe a moment when you felt deeply inspired by an artistic or aesthetic experience?” This encourages detailed, contextual responses beyond a simple “yes” or “no.”
3.1 PSI Questions Design
How can we design questions more effectively and obtain textual information with greater depth in measuring personality? McAdams’ three levels of personality theory and the corresponding, theory-informed interviews offer a valuable direction (e.g., McAdams, 1995, 1996, 2001).
According to McAdams’s theory, personality can be divided into three levels: traits (broad personality characteristics), personal concerns (goals, values, strategies), and narrative identity (the story individuals tell themselves about their lives). The latter two levels, in particular, reveal how individuals act in specific situations and developmental stages and how they organize and make sense of their life experiences.
The key to obtaining more meaningful textual information lies in guiding individuals toward higher levels of self-expression. Rather than simply recording mundane daily events, we should encourage reflection on pivotal life moments, significant relationships, and future aspirations. For example, effective questions could be: “Can you describe an event that changed the trajectory of your life?” or “Tell me about a moment you are most proud of.” By designing questions like these, we can elicit deeper, more meaningful narratives, enriching the personality information.
We adapted and modified McAdams (1995, 1996, 2001) theory-based interview and narrative identity approach, incorporating elements from the structured interview of the FFM (SIFFM; Trull et al., 1998). The initial draft of the question pool was prepared and revised by subject matter experts (SMEs) through discussion. The SMEs included two doctoral students, one postdoctoral researcher, and one professor, all specializing in personality psychology. Subsequently, we conducted pilot testing with six undergraduate research assistants from a personality research lab.
As a result, the PSI, comprising 32 questions, was developed, as shown in Table 3 in Appendix B.
3.2 PSI Dataset
Using this set of questions, we designed a chatbot to conduct structured interviews, facilitating data collection (Institutional Review Board approval was obtained). Because data collection is ongoing, we used responses collected through the end of December 2024. After excluding incomplete responses and those failing attention checks, the final sample consisted of 357 participants. On average, participants were in their early 30s (M = 33.30 years, SD = 13.06), with 52.40% identifying as men, and 44.80% as women.
@@ 3.2 PSI Dataset
To ensure a demographically diverse and broadly representative sample, we recruited participants from both undergraduate and working adult populations, ensuring a diverse range of backgrounds in terms of educational, employment, and life experience. Undergraduate participants were recruited from a large public university in the Midwest and received research credit for their participation. Working adult participants were sourced from two widely recognized, high-quality crowdsourcing platforms—Prolific [^2] https://www.prolific.com/ and CloudResearch Connect [^3] https://connect.cloudresearch.com/—which are known for their diverse and demographically representative participant pools. All working adult participants were compensated for their time. The average interview duration was 34 minutes.
In addition to completing the structured interview, participants provided demographic information and responded to both the personality scale and the personality-related behaviors scale (see Appendix D for detailed descriptions of data collection and examples of the dataset). Responses to PSI were subsequently incorporated into the prompt to generate simulated data (see Appendix F for the specific prompts used in the simulation process).
4 Experimental Setup
We designed three primary experiments to verify the effectiveness of the PSI method and its advantages over other methods: (1) Response Similarity (§4.2): evaluating the degree of similarity between responses generated by LLMs based on the PSI method and actual corresponding human responses on personality scales; (2) Human Personality Distribution Simulation (§4.3): comparing the PSI method with other methods in simulating human personality distributions; (3) Personality-Related Behavioral Performance (§4.4): testing the models’ performance using the PSI method to simulate personality-related behaviors. General settings shared by all experiments are described in §4.1, while specific differences are detailed in their respective sections.

### v2

# Generative Personality Simulation via Theory-Informed Structured Interview

@@ Abstract
Despite their potential as human proxies, LLMs often fail to generate heterogeneous data with human-like diversity, thereby diminishing their value in advancing social science research. To address this gap, we propose a novel method to incorporate psychological insights into LLM simulation through the Personality Structured Interview (PSI [^1] Code and data: https://github.com/isle-dev/PSI.). PSI leverages psychometric scale-development procedures to capture personality-related linguistic information from a formal psychological perspective. To systematically evaluate simulation fidelity, we developed a measurement theory grounded evaluation procedure that considers the latent construct nature of personality and evaluates its reliability, structural validity, and external validity. Results from three experiments demonstrate that PSI effectively improves human-like heterogeneity in LLM-simulated personality data and predicts personality-related behavioral outcomes. We further offer a theoretical framework for designing theory-informed structured interviews to enhance the reliability and effectiveness of LLMs in simulating human-like data for broader psychometric research.
Generative Personality Simulation via Theory-Informed Structured Interview
Pengda Wang1, Huiqi Zou3, Han Jiang4, Hanjie Chen2, Tianjun Sun1, Xiaoyuan Yi5, Ziang Xiao4, and Frederick L. Oswald1 1Department of Psychological Sciences & 2Department of Computer Science, Rice University 3Department of Electrical and Computer Engineering, Northeastern University 4Department of Computer Science, Johns Hopkins University 5Microsoft Research Asia {pw32,ts110,fo3}@rice.edu; {ziang.xiao}@jhu.edu
@@ 1 Introduction
The discipline of personality psychology seeks to understand how individual differences shape significant life outcomes and trajectories. Decades of empirical work have demonstrated that personality traits predict a wide range of significant domains, including career success, mental and physical health, interpersonal relationships, and overall well-being (e.g., Judge et al., 2002; Roberts et al., 2007; Robins et al., 2002). Therefore, personality assessment has often been used to inform personnel selection and clinical interventions. Moreover, the relevance of personality research is rapidly expanding into the field of Artificial Intelligence (AI), where it informs the design of more adaptive, human-centered, and personalized systems.
@@
Figure 1: Overview of the development of the personality structured interview and experimental implementation.
However, studying personality requires large amounts of psychological data, such as personality trait scores and behavioral indicators. These data collection demands pose substantial logistical and financial challenges for researchers. Moreover, certain types of data are particularly difficult to obtain. For example, studies focusing on individuals with elevated levels of psychopathy or narcissism often struggle to obtain adequate sample sizes due to the rarity and sensitivity of these traits (e.g., Lynam and Widiger, 2001). Other research questions require longitudinal designs to capture personality development or dynamic patterns of social interaction over time, further compounding the complexity and resource demands (e.g., Damian et al., 2019; Roberts et al., 2006). If a method could accurately simulate human personality distributions, it would accelerate personality research by offering a more scalable and cost-effective way to support data collection and experimentation (Messeri and Crockett, 2024).
In this context, the current work investigates the potential of LLMs to simulate human-like responses in psychometric research. Although prior studies have shown such promise, indicating that LLMs can generate responses for personality scales that reflect personality traits resembling those of humans (e.g., Huang et al., 2023, Lee et al., 2024), those approaches exhibit notable limitations, particularly in their ability to capture individual differences at the item-level and to reproduce the heterogeneity observed in human responses (e.g., Wang et al., 2024). Item-level data provides a more granular understanding of how traits manifest in specific individuals rather than relying solely on broad trait averages. Furthermore, modeling response heterogeneity is crucial for reflecting the variability and complexity of real-world human behavior. Additionally, current evaluation practices are often misaligned with the standards of psychological research. For example, prior research has predominantly assessed personality simulation data at the aggregated level, neglecting item- and facet-level analyses as well as the psychometric properties of the data. The evaluation gap undermines psychologists’ confidence in the validity of such methods.
To address these research gaps, we investigated the use of Personality Structured Interview (PSI) transcripts as a means of guiding LLMs in simulating personality data (see Figure 1). The PSI comprises theory-driven questions paired with responses, designed to elicit narrative-based information pertinent to the personality constructs targeted in LLM simulations. This theory-informed approach enables the expression of personality in a manner that enhances the human-like heterogeneity and representativeness of the simulated data. To bridge the evaluation gap, we developed a comprehensive framework for assessing personality simulation data grounded in measurement theory (e.g., Cronbach, 1951; Cronbach and Meehl, 1955). This evaluation framework explicitly accounts for the hierarchical structure of personality measurement, wherein observed responses at the item-level map onto latent traits at the domain-level. It further supports a wide range of analytic strategies, from basic descriptive statistics to more advanced assessments of psychometric validity and reliability.
With this framework, we conducted three experiments to evaluate the efficacy of the PSI method in: (1) replicating individual-level personality data, (2) simulating distributions of personality approximate those observed in human populations, and (3) capturing personality-related behaviors aligned with established findings. Our findings demonstrate that the PSI approach enhances human-like variability of simulated personality data and effectively captures the patterns of personality-relevant behavior.
In summary, our paper makes three key contributions: (1) a theory-informed LLM-based simulation method, PSI, for personality research, along with a framework for developing such interview protocols; (2) the release of a dataset comprising 357 structured interview transcripts; and (3) an evaluation framework for LLM-simulated psychometric data, grounded in measurement theory. We further discuss the potential of the PSI method and the dataset to advance AI and psychology research (see Appendix G). The interview question development framework underlying PSI can be generalized to capture information relevant to other constructs (e.g., value), thereby facilitating the development of more theory-aligned data simulations and enhancing the alignment of LLMs with humans.
2 Related Works
Personality Definition and Construct
@@ 2 Related Works
MacKinnon (1944) proposed two complementary definitions of personality. One emphasizes internal factors like temperament and interpersonal strategies that drive consistent behavior across time, situations, and cultures. The other focuses on interpersonal characteristics as perceived by others, linking personality to reputation. The former highlights internal drives, while the latter centers on external behaviors and social perception. Together, these perspectives underscore personality’s role in shaping thought patterns and behaviors in social interactions (e.g., Hogan et al., 1996).
@@
Personality encodes rich and complex information in language and text (e.g., Goldberg, 1990; Saucier and Goldberg, 2001). The Five Factor Model (FFM) of personality is extensively researched (e.g., Costa and McCrae, 2008; John, 1999; McCrae and Costa Jr, 1997); it is based on the lexical hypothesis, which posits that individual differences that are important in human interactions (e.g., have survival value across cultures) tend to become encoded in language. The five factors are Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (OCEAN). This theoretical foundation has evolved over decades, from Galton (1884)’s early work on trait descriptors, through Allport and Odbert (1936)’s lexical studies, and into more systematic factor-analytic approaches by scholars such as Norman (1963) and Goldberg (1990). These developments have shaped our modern understanding of personality structure (see Appendix F for further details). Given the linguistic basis of personality, LLMs trained on vast and diverse text corpora are well-equipped to model personality data from natural language use.
Generative Personality Simulation
@@
Current approaches to simulating personality data with LLMs exhibit several notable limitations. The Persona-Chat dataset, introduced by Zhang et al. (2018), was originally designed to enhance chit-chat models through increased personalization. As such, its primary goal is to foster conversational engagement rather than to accurately model underlying personality constructs. Another approach, the adjective-based categorization method (Shape method) proposed by Serapio-García et al. (2023), enables the simulation of profile-specific patterns by targeting high or low expressions of individual personality dimensions. However, this method struggles to realistically capture personality distributions, as it constrains each case to a single trait dimension—oversimplifying the inherently multidimensional and hierarchical nature of personality structure (e.g., Kachur et al., 2020). To address these limitations, we explore the use of PSI as a means to generate more nuanced, ecologically valid, and representative simulations of personality data.
3 Personality Structured Interview for LLM Simulation
The PSI differs from earlier interview datasets, such as the Life Interview transcripts (Park et al., 2024), which ask broad, general questions about a person’s life. In contrast, the PSI is purposefully designed to elicit linguistic cues associated with personality constructs. Based on psychometric scale development methods, we developed a structured framework to guide the formulation of interview questions (see Appendix G).
Simply rephrasing scale items into open-ended questions often fails to elicit meaningful elaboration. Converting an item ‘‘Values artistic, aesthetic experiences’’ into ‘‘Do you value artistic, aesthetic experiences?’’ typically yields a brief ‘‘yes’’ or ‘‘no’’ answer, offering little insight (Trull et al., 1998). Such questions also diverge from the behavioral prompts commonly used in structured interviews, which are more effective for eliciting rich, narrative data (e.g., Campion et al., 1997). To explore aesthetic values, asking ‘‘Describe a moment when you felt inspired by an artistic or aesthetic experience’’ invites deeper, more contextual responses.
3.1 Personality Structured Interview Development Process
The overall process of developing a structured interview designed to assess the target construct is detailed in Appendix G. Here, let $\mathcal{C}$ denote the personality construct of interest. We present a complete example, formulated as a functional chain, to illustrate how the final set of structured interview questions, $\mathcal{Q}_{\text{final}}$, is produced: $\mathcal{C}\xrightarrow{f_{1}}\mathcal{I}\xrightarrow{f_{2}}\mathcal{B}\xrightarrow{f_{3}}\mathcal{Q}_{1}\xrightarrow{f_{4}}\mathcal{Q}_{2}^{(0)}\xrightarrow{f_{5}~(\text{iterative})}(\mathcal{R},\mathcal{V})\xrightarrow{f_{6}}\mathcal{Q}_{\text{final}}$.
In the original steps, we first need to identify the behavior/perception indicators $\mathcal{I}$ demonstrated by $\mathcal{C}$ ($\mathcal{C}\xrightarrow{f_{1}}\mathcal{I}$), and based on $\mathcal{I}$ to build the initial blueprint $\mathcal{B}$ for the structured interview questions ($\mathcal{I}\xrightarrow{f_{2}}\mathcal{B}$). Then, based on $\mathcal{B}$, we can generate the initial question pool $\mathcal{Q}_{1}$ ($\mathcal{B}\xrightarrow{f_{3}}\mathcal{Q}_{1}$). However, given the substantial body of research on personality assessment and constructs—such as narrative identity theory (McAdams, 1995, 1996, 2001) and the Structured Interview for the Five-Factor Model (SIFFM; Trull et al., 1998) [^2] McAdams’ three levels of personality (traits, personal concerns, and narrative identity) as well as other established personality structured interviews, highlight the importance of thoughtfully guiding individuals toward deeper self-expression when designing such interviews. Instead of focusing on routine events or superficial details, we should prompt reflection on pivotal moments, meaningful relationships, and future aspirations. For example, questions like ‘‘Can you describe an event that changed the trajectory of your life?’’ or ‘‘Tell me about a moment you are most proud of.’’ can elicit rich, narrative-driven insights into personality.—four subject matter experts ($\text{SMEs}=\{\text{Ph.D.}_{student1},\text{Ph.D.}_{student2},\text{Postdoc},\text{Professor}\}$), all specializing in Personality Psychology, adapted and modified these existing theory and interview questions accordingly to generate the initial question pool: $\mathcal{Q}_{1}$ ($\mathcal{Q}_{1}=\text{adapt}(\mathcal{Q}_{\text{prior}},\text{SMEs})$).
Following this, we conducted pilot testing with six undergraduate research assistants affiliated with a personality research laboratory ($\mathcal{D}_{\text{pilot}}$). Based on their feedback, the interview questions were further revised by SMEs ($\mathcal{Q}_{2}=f_{4}(\mathcal{Q}_{1},\text{Stat}(\mathcal{D}_{\text{pilot}}))$). Then we conducted field testing with a convenience sample of university participants ($\mathcal{D}_{\text{field}}$). During field testing, the other three doctoral students in Industrial/Organizational and Personality Psychology served as independent observers: they reviewed the interview transcripts and provided observer ratings of participants’ personalities to validate the reliability ($\mathcal{R}$) and validity ($\mathcal{V}$) of the structured personality interview questions ($\mathcal{R}^{(t)},\mathcal{V}^{(t)}=f_{5}(\mathcal{Q}_{2}^{(t)},\mathcal{D}_{\text{field}})$). This is an iterative process—if the evaluated metrics do not meet the predefined thresholds $\tau_{\text{rel}}$ and $\tau_{\text{val}}$, the questions are further revised ($\mathcal{Q}_{2}^{(t+1)}=f_{\text{revise}}(\mathcal{Q}_{2}^{(t)},\mathcal{R}^{(t)},\mathcal{V}^{(t)})$); otherwise, the process produces $\mathcal{Q}_{\text{final}}$ ($\mathcal{R},\mathcal{V}\xrightarrow{f_{6}}\mathcal{Q}_{\text{final}}$).
Our field testing results have shown that the average correlation between observer ratings and self-report ratings ($\mathcal{V}$) was 0.36, and the average inter-rater reliability (intraclass correlation; $\mathcal{R}$) was 0.76. These values are consistent with theoretical expectations—meta-analytic research suggests an average self–other rating correlation of 0.36 ($\tau_{\text{val}}$) for the Big Five traits (e.g., Connolly et al., 2007), and the inter-rater reliability obtained exceeds the commonly accepted threshold of 0.70 ($\tau_{\text{rel}}$). Detailed results from the validation phase are presented in Appendix B. As a result, $\mathcal{Q}_{\text{final}}$, comprising 32 carefully developed questions, was developed, as shown in Table 7 in Appendix B.
3.2 Personality Structured Interview Dataset
@@
Data were collected online using the finalized set of interview questions, $\mathcal{Q}_{\text{final}}$ (Institutional Review Board approval was obtained; protocol number and institution are masked for blind review). Participants were sampled from both undergraduate and working adult populations. Undergraduate participants were recruited from a large public university in the U.S. Midwest and received course credit for their participation. Working adults were recruited via two widely recognized, high-quality crowdsourcing platforms—Prolific [^3] https://www.prolific.com/ and CloudResearch Connect [^4] https://connect.cloudresearch.com/—both known for providing access to diverse and demographically representative samples. All working adult participants were compensated for their time. On average, each structured interview session lasted approximately 34 minutes.
In addition to responding to $\mathcal{Q}{\text{final}}$, participants provided demographic information and completed a standardized personality inventory (BFI-2; Soto and John, 2017b), along with scales measuring personality-relevant behaviors (organizational citizenship behavior & counterproductive work behavior; Spector and Fox, 2010) (see Appendix C for detailed descriptions and examples). Their responses to $\mathcal{Q}{\text{final}}$ were subsequently incorporated into prompt templates used for generating simulated data (see Appendix D for the specific simulation prompts). After excluding incomplete responses and those that failed attention checks, the final analytic sample included 357 participants. The average age of the participants was 33.30 years, with a standard deviation of 13.06 years. In terms of gender identity, 52.40% identified as male and 44.80% as female. The racial composition of the sample aligned with national demographic trends, with 70.59% identifying as White, 10.92% as Black or African American, 9.24% as Hispanic/Latinx, and 7.84% as Asian, and 1.40% as Other.
4 Experiments and Results